# How to Securely Shuffle? A survey about Secure Shufflers for privacy-preserving computations

### v1

@@ Abstract.
This work was supported by the Netherlands Organization for Scientific Research (NWO) in the context of the SHARE project [CS.011], the ANR project ANR-20-CE23-0013 PMR, and by the Horizon Europe project HE TRUMPET
journal: CSURccs: Security and privacyccs: Computing methodologies Machine learning
@@ 1. Introduction
This line of work originates from Ishai et al. (2006) that proposed, in 2006, secure shufflers to perform privacy-perserving data aggregation. Recent works in Differential Privacy (DP) (Bittau et al., 2017; Balle et al., 2019a, b; Cheu et al., 2019; Erlingsson et al., 2019; Cheu, 2021) revived this interest in secure shufflers by formalizing the “privacy amplification” they provide. Based on DP analysis, shuffling-based private computations offer privacy-utility trade-off close to private computations performed by a trusted third party (i.e., a trusted party offering the optimal trade-off in DP). Thus, secure shufflers have been presented as an efficient solution to build privacy-preserving computations without requiring a trusted party.
@@
In addition to the privacy amplification in DP, secure shufflers have found applications in numerous fields from privacy-preserving software monitoring (Bittau et al., 2017) to Federated Learning (FL) (Konečný et al., 2016, 2017); a decentralized ML paradigm in which data owners collaboratively without sharing their personal data. In these applications, the shufflers enable to compute efficiently and privately a form of data aggregation without revealing the individual values. Recently, Wang et al. (Wang et al., 2025) broadened the scope of applications as they highlighted the potential of secure shufflers in “private individual computations” (e.g., taxi hailing).
@@
Most papers introducing such shuffling-based computations (e.g., (Ghazi et al., 2019; Cheu et al., 2019)) use an abstract shuffler rather than a concrete implementation. An advantage of considering an abstract shuffler is to focus on the statistical properties, which are relevant for the applications, and plug in later a concrete shuffler implementation. However, not all secure shufflers are equally appropriate for all applications, as shuffler and application papers make their own explicit or implicit security assumptions, which must be compatible.
@@
Understanding properties of secure shuffler implementations has several benefits. First, it allows for a fairer comparison between private computation algorithms. For example, existing papers do not tell whether shuffling-based data aggregation is more scalable than masking-based data aggregation (Mansouri et al., 2023), since most papers on shuffling-based computations do not detail their shuffler implementation. Second, it contributes to a better formalization of the concept of “secure shuffler”. The literature relies on a vague/intuitive definition leaving some blind spots; e.g., several papers (Ghazi et al., 2019; Cheu et al., 2019; Kairouz et al., 2021) present mix networks as an example of secure shufflers, without mentioning the possible de-anonymization and censorship attacks (Shmatikov and Wang, 2006; Sun et al., 2015) Third, it could lead to more consensus on what technique is appropriate in which use cases. For example, Ghazi et al. (Ghazi et al., 2019) present mix nets as “highly scalable” while Gordon et al. (Gordon et al., 2022) point at the same shufflers as too expensive, which may seem contradictory without a deeper understanding of the context.
@@
Implementing an ideal shuffler (as defined in the theory) is not straightforward. Bagdasaryan et al. (2022) pointed out the need for a thorough understanding of secure shuffler implementations as they highlighted that only one DP paper (Bittau et al., 2017) detailed their implementation. Our survey comes to fill this gap identified by (Bagdasaryan et al., 2022).
@@
    We identify and compare 26 existing secure protocols fulfilling the minimum requirements of a secure shuffler. We adapt their original security definitions to a uniform set of security properties that supports our comparison.
@@
    We survey the applications of secure shufflers in privacy-preserving computations.
@@
    We identify several open problems (e.g., robustness to data poisoning).
@@
    Guidelines to choose a shuffler based on the properties of a privacy-preserving application.
@@ 2.2. Differential Privacy
Dwork (Dwork, 2006) introduced differential privacy (DP) to characterize the privacy leakage of any computation with public output:
@@
The space $\mathbb{X}$ corresponds to the secret value space (i.e., each data owner $P_{i}$ owns a private value $x_{i}\in\mathbb{X}$). The parameter $\epsilon\in[0,\infty)$ controls the privacy loss: smaller $\epsilon$ implies more privacy. To ensure strong privacy, $\epsilon$ is usually chosen to be below 1 (Abadi et al., 2016). The parameter $\delta\in[0,1)$ relaxes the definition and allows for rare cases where the privacy loss would be higher than $\epsilon$. To ensure strong privacy, the probability $\delta$ should be below $1/|D|$ (with $D\in\mathbb{X}^{N}$, a dataset) (Abadi et al., 2016).
@@
Several papers have proposed strategies to “amplify the privacy”; i.e., reducing the amount of noise needed to reach a certain privacy level. These solutions include iterative protocols (Feldman et al., 2018), decentralization (Cyffers and Bellet, 2021), random check-ins(Balle et al., 2020b) and shuffling (Erlingsson et al., 2019) (detailed in Subsection 5.2).
@@ 2.3.2. Secret Sharing
Shares are a set of values that, once gathered, reveal a secret value. Individually, the shares leak no information about the secret value. A secret value $x$ is considered shared when several parties own a share; $P_{j}$ owning the share $x_{j}$. Several secret sharing schemes exist, e.g., additive shares are numbers drawn randomly subject to the constraint that their sum (in a finite field) reveals the shared value: $\sum_{j}x_{j}=x$. Recent works also enable more complex forms of secret sharing, such protocols to share a function (Boyle et al., 2015).
@@ 2.3.3. Multi-party computation (MPC)
MPC (Evans et al., 2018) is a set of techniques enabling several parties to compute a function over secret inputs. Many MPC protocols rely notably on secret sharing schemes. The main bottleneck of MPC is the number of parties: while they provide high security, state-of-the-art protocols scale quadratically with the number of shareholders (i.e., they would not be applied for more than a handful of parties).
@@ 2.3.4. Zero-knowledge proofs (ZKP)
ZKPs (Goldreich and Oren, 1994) are techniques to prove properties about ciphertexts (e.g., that two encrypted values are equal or that an encrypted value is in a given range) without disclosing anything else. A ZKP involves two agents: the prover (who provides the proof) and the verifier (who verifies its correctness).
@@ 2.3.5. Trusted Execution Environment (TEE)
A trusted execution environment (TEE) (Sabt et al., 2015) (sometimes called “secure enclave”) is a secure environment in which confidentiality and integrity for the processed data and code are ensured. Even the hardware owner cannot read or modify the data processed in the TEE. Anyone can audit the code executed in the environment, and the hardware design ensures that the data processed inside remains confidential.
They require no trust other than in the manufacturer. Many major chip manufacturers have already designed such environments; e.g., Intel SGX (Intel, 2015).
@@ 2.4. Basic shuffling-based data aggregation
Ishai et al. (Ishai et al., 2006) introduced the first shuffling-based private computation. In particular, they presented a shuffling-based protocol to aggregate private values. This work is the seminal work that inspired most recent works on shuffling-based private computations (Balle et al., 2019a; Wang et al., 2025).
@@
Ishai et al. (2006) showed that each data owner only needs to send a logarithmic number of shares in order to have shares indistinguishable from perfect randomness. To reconstruct the input secret values based on the shuffled random shares, an adversary would need to compute an exponential number of share combinations. The shuffling combined with the share randomization then guarantees that an attacker cannot realistically harm the input privacy, but allows to efficiently perform a data aggregation while preserving input privacy.
@@ 3.1. Threat models
Threat models characterize agents or behaviors that threaten a desirable security property of a protocol (e.g., correctness or privacy). When defining a secure protocol, it is necessary to specify a threat model against which the security properties are guaranteed. In other words, the desirable property of a protocol holds as long as no adversary is more powerful (i.e., has more knowledge or computational power) than what is defined in the threat model. This subsection reminds the terminology commonly used in security literature (Evans et al., 2018).
@@
Secure protocols typically guarantee data privacy against a proportion of malicious/semi-honest agents. For instance, Clarion (Eskandarian and Boneh, 2022) is secure as long as one out of $M$ servers (involved in the protocol) is honest. On the other hand, RPM (Lu and Kate, 2023) is secure as long as there is a majority of honest servers involved in the protocol.
@@
While most secure protocols restrict their threat model to the agents actively involved in the protocol, anonymous communication protocols can also be the target of traffic analysis attacks (Murdoch and Danezis, 2005). Some works then include in their threat model a “global adversary” able to capture the network traffic of the agents. A “global adversary” is typically a state or an internet service provider.
@@ 3.2. Shuffler definition
Our starting point is the definition of a shuffler $\mathcal{S}$. Based on papers studying shuffling-based DP (Cheu et al., 2019; Cheu, 2021), we propose the following naive definition:
@@
Figure 1. Abstract shuffler design: (1) The senders pre-process their messages using $\mathcal{P}$, (2) They send their preprocessed messages to the mixer $\mathcal{M}$, (3) The mixer outputs a permutation of the messages, and (4) an analyser can use the shuffled messages in a privacy-preserving computation.
[…]
[Schema of an abstract shuffler]Schema of an abstract shuffler: three data owners preprocess their private data and send them to a mixer M; transmitting the shuffled output to an analyzer.
@@
This informal definition fits what the literature (Kairouz et al., 2021) implicitly considers as a “secure shuffler” (also referred to as “oblivious shuffler” in (Bittau et al., 2017; Gordon et al., 2022)). However, this definition is imprecise. Thus, we propose to formalize this notion using four new security properties to characterize a shuffler $\mathcal{S}$: Anonymity, Correctness, Sender disruption resistance, Server disruption resistance.
@@ 3.3.1. Anonymity
This is the minimum requirement for shuffling-based private computations. More precisely, one is usually interested in sender anonymity because only the origin of the messages must remain anonymous. It is even the core concern of shuffler application paper that give more importance to the concept of “anonymous messaging” than to shuffling itself (Ishai et al., 2006; Cheu et al., 2019; Ghazi et al., 2020a).
However, anonymity is never formally defined by these papers. For example, Ishai et al. (2006) states that the shuffling-based private aggregation requires a “sufficient uncertainty.”
@@
While the first anonymization techniques (Chaum, 1981, 1988) appeared in the 1980s, the first anonymity formalization paper (Feigenbaum et al., 2007) was only published in 2007. Thus, most anonymization systems (up until now) have been built upon informal anonymity definitions. For example, even recent constructions such as Clarion (Eskandarian and Boneh, 2022) prove the security of their MPC protocol, but do not prove the anonymity itself; (Eskandarian and Boneh, 2022) does not even cite any formalization framework.
Pfitzmann and Hansen (2010) proposed a recognized terminology for all the words related to anonymity (Lu et al., 2019a). These definitions are informal but are a commonly accepted starting point. They give the following definition for sender anonymity:
@@
Finding a formal definition for anonymity is a complex problem because anonymity is context-dependent. Several papers (Backes et al., 2013; Gelernter and Herzberg, 2013; Kuhn et al., 2019) proposed formal anonymity definitions, but they remain imperfect. For example, some of them exclude some de-anonymization attacks (e.g., traffic analysis attacks in (Backes et al., 2013)), which creates a blind spot in the theory that struggles to fit the reality of anonymity. Furthermore, Lu et al. (2019a) surveyed these works on anonymity formalizations and concluded that the formalization proposed “is still a theoretic approach and lacks a practical verification.” Hence, anonymity formalization is still an open problem for the security community that is still looking for a standard and general solution.
@@
This definition means that an adversary can distinguish honest senders only with negligible probability. Suppose all the messages are unique, the probability of $m$ being sent by the honest party $P$ (i.e., $\Pr[m\text{ sent by }P|\mathcal{V}(A)]$) has two possible values: 00 if the adversary controls the sender of $m$, and $\frac{1}{\lvert\mathcal{H}\rvert}$, otherwise. If the messages are not all unique, we can have $q$ honest senders who sent the same message $m$; the probability of a (honest) $P$ being the sender of $m$ is $\frac{q}{\lvert\mathcal{H}\rvert}$.
@@
Our survey focuses on sender anonymity as it is the form of anonymity required by secure shufflers. However, some anonymous communication systems provide other forms of anonymity such as one-to-one anonymous communications (van den Hooff et al., 2015; Eskandarian et al., 2021) (i.e., the two communicating agents know the identity of their correspondent, but there is a guarantee that no third party knows these agents are communicating). This is useful in a number of applications but incompatible with sender anonymity since the receiver knows the sender’s identity. Hence, such systems cannot be used as secure shufflers. Among all anonymous communication systems, only “anonymous broadcasters” satisfy the secure shuffling needs.
Over the years, several works including (Edman and Yener, 2009; Sasy and Goldberg, 2023; Shirali et al., 2022; Shirazi et al., 2018) have surveyed anonymous communication systems As Sasy and Goldberg (2023) is the most recent and complete survey, we will focus our discussion on it, but our remarks hold for all surveys on anonymous communications. While both (Sasy and Goldberg, 2023) and our work gravitate around the notion of anonymity, they provide different perspectives because they do not focus on the same functionality. On the one hand, they focus on anonymous communications. On the other hand, we focus on secure shuffling.
Since we focus on different functionalities, our work relies on a different set of definitions. For example, the concept of correctness (defined in the next subsubsection) is not studied in (Sasy and Goldberg, 2023). Indeed, correctness is an important notion in privacy-preserving computations as we want to avoid a bias in the output results. The anonymous communications literature is much more interested in the notion of censorship resistance (which is implied by correctness but is not equivalent).
This scope difference has a visible consequence: while some solutions are covered in both works, some solutions are only present in (Sasy and Goldberg, 2023) (e.g., one-to-one anonymous communications), and others are only present in our work (e.g., TEE-based shufflers).
A core contribution of our work is to identify secure protocols able to implement a secure shuffler, even though these protocols were not initially designed for it. While some anonymous communications systems can implement this functionality, many other secure protocols provide the sender anonymity necessary to secure shuffling. For example, Section 4 highlights ZKP protocols initially used for electronic voting (Adida, 2008) and TEE solutions imagined for secure contact discovery (Ngai et al., 2024).

### v2

@@ Abstract.
journal: CSURccs: Security and privacyccs: Security and privacy Cryptographyccs: Computing methodologies Machine learning
@@ 1. Introduction
This line of work originates from ishai_cryptography_2006 that proposed, in 2006, secure shufflers to perform privacy-preserving data aggregation. Recent works in Differential Privacy (DP) (bittau_prochlo_2017; balle_improved_2019; balle_privacy_2019; cheu_distributed_2019; erlingsson_amplification_2019; cheu_differential_2021) revived this interest in secure shufflers by formalizing the “privacy amplification” they provide. Based on DP analysis, shuffling-based private computations offer privacy-utility trade-off close to private computations performed by a trusted third party (i.e., a trusted party offering the optimal trade-off in DP). Thus, secure shufflers have been presented as an efficient solution to build privacy-preserving computations without requiring a trusted party.
@@
In addition to the privacy amplification in DP, secure shufflers have found applications in numerous fields from privacy-preserving software monitoring (bittau_prochlo_2017) to Federated Learning (FL) (konecny_federated_2016; konecny_federated_2017); a decentralized ML paradigm in which data owners collaboratively without sharing their personal data. In these applications, the shufflers enable to compute efficiently and privately a form of data aggregation without revealing the individual values. Recently, Wang et al. (wang_beyond_2025) broadened the scope of applications as they highlighted the potential of secure shufflers in “private individual computations” (e.g., taxi hailing).
@@
Most papers introducing such shuffling-based computations (e.g., (ghazi_scalable_2019; cheu_distributed_2019)) use an abstract shuffler rather than a concrete implementation. An advantage of considering an abstract shuffler is to focus on the statistical properties, which are relevant for the applications, and plug in later a concrete shuffler implementation. However, not all secure shufflers are equally appropriate for all applications, as shuffler and application papers make their own explicit or implicit security assumptions, which must be compatible.
@@
Understanding properties of secure shuffler implementations has several benefits. First, it allows for a fairer comparison between private computation algorithms. For example, existing papers do not tell whether shuffling-based data aggregation is more scalable than masking-based data aggregation (mansouri_sok_2023), since most papers on shuffling-based computations do not detail their shuffler implementation. Second, it contributes to a better formalization of the concept of “secure shuffler”. The literature relies on a vague definition leaving some blind spots; e.g., several papers (ghazi_scalable_2019; cheu_distributed_2019; kairouz_advances_2021) present mix networks as an example of secure shufflers, without mentioning the possible de-anonymization and censorship attacks (shmatikov_timing_2006; sun_raptor_2015). Third, it could lead to more consensus on what technique is appropriate in which use cases. For example, Ghazi et al. (ghazi_scalable_2019) present mix nets as “highly scalable” while Gordon et al. (gordon_spreading_2022) point at the same shufflers as too expensive, which may seem contradictory without a deeper understanding of the context.
@@
Implementing an ideal shuffler (as defined in the theory) is not straightforward. bagdasaryan_towards_2022 pointed out the need for a thorough understanding of secure shuffler implementations as they highlighted that only one DP paper (bittau_prochlo_2017) detailed their implementation. Our survey comes to fill this gap.
@@
    Identify and compare 26 existing secure protocols fulfilling the minimum requirements of a secure shuffler.
@@
    Adapt their original security definitions to a uniform set of security properties that supports our comparison.
@@
    Survey the applications of secure shufflers in privacy-preserving computations.
@@
    Identify several open problems (e.g., robustness to data poisoning).
  5. (5)
    Introduce guidelines to choose a shuffler based on the properties of a privacy-preserving application.
@@ 2.2. Differential Privacy
Dwork (dwork_differential_2006) introduced differential privacy (DP) to characterize the privacy leakage of any computation with public output:
@@
The space $\mathbb{X}$ corresponds to the secret value space (i.e., each data owner $P_{i}$ owns a private value $x_{i}\in\mathbb{X}$). The parameter $\epsilon\in[0,\infty)$ controls the privacy loss: smaller $\epsilon$ implies more privacy. To ensure strong privacy, $\epsilon$ is usually chosen to be below 1 (abadi_deep_2016). The parameter $\delta\in[0,1)$ relaxes the definition and allows for rare cases where the privacy loss would be higher than $\epsilon$. To ensure strong privacy, the probability $\delta$ should be below $1/|D|$ (with $D\in\mathbb{X}^{N}$, a dataset) (abadi_deep_2016).
@@
Several papers have proposed strategies to “amplify the privacy”; i.e., reducing the amount of noise needed to reach a certain privacy level. These solutions include iterative protocols (feldman_privacy_2018), decentralization (cyffers_privacy_2021), random check-ins(balle_privacy_2020) and shuffling (erlingsson_amplification_2019) (detailed in Subsection 5.2).
@@ 2.3.2. Secret Sharing
Shares are a set of values that, once gathered, reveal a secret value. Individually, the shares leak no information about the secret value. A secret value $x$ is considered shared when several parties own a share; $P_{j}$ owning the share $x_{j}$. Several secret sharing schemes exist, e.g., additive shares are numbers drawn randomly subject to the constraint that their sum (in a finite field) reveals the shared value: $\sum_{j}x_{j}=x$. Recent works also enable more complex forms of secret sharing, such protocols to share a function (boyle_function_2015).
@@ 2.3.3. Multi-party computation (MPC)
MPC (evans_pragmatic_2018) is a set of techniques enabling several parties to compute a function over secret inputs. Many MPC protocols rely notably on secret sharing schemes. The main bottleneck of MPC is the number of parties: while they provide high security, state-of-the-art protocols scale quadratically with the number of shareholders (i.e., they would not be applied for more than a handful of parties).
@@ 2.3.4. Zero-knowledge proofs (ZKP)
ZKPs (goldreich_definitions_1994) are techniques to prove properties about ciphertexts (e.g., that two encrypted values are equal or that an encrypted value is in a given range) without disclosing anything else. A ZKP involves two agents: the prover (who provides the proof) and the verifier (who verifies its correctness).
@@ 2.3.5. Trusted Execution Environment (TEE)
A trusted execution environment (TEE) (sabt_trusted_2015) (sometimes called “secure enclave”) is a secure environment in which confidentiality and integrity for the processed data and code are ensured. Even the hardware owner cannot read or modify the data processed in the TEE. Anyone can audit the code executed in the environment, and the hardware design ensures that the data processed inside remains confidential.
They require no trust other than in the manufacturer. Many major chip manufacturers have already designed such environments; e.g., Intel SGX (intel_software_2015).
@@ 2.4. Basic shuffling-based data aggregation
Ishai et al. (ishai_cryptography_2006) introduced the first shuffling-based private computation. In particular, they presented a shuffling-based protocol to aggregate private values. Their work is the seminal work that inspired most recent works on shuffling-based private computations (balle_improved_2019; wang_beyond_2025).
@@
ishai_cryptography_2006 showed that each data owner only needs to send a logarithmic number of shares in order to have shares indistinguishable from perfect randomness. To reconstruct the input secret values based on the shuffled random shares, an adversary would need to compute an exponential number of share combinations. The shuffling combined with the share randomization then guarantees that an attacker cannot realistically harm the input privacy, but allows to efficiently perform a data aggregation while preserving input privacy.
@@ 3.1. Threat models
Threat models characterize agents or behaviors that threaten a desirable security property of a protocol (e.g., correctness or privacy). When defining a secure protocol, it is necessary to specify a threat model against which the security properties are guaranteed. In other words, the desirable property of a protocol holds as long as no adversary is more powerful (i.e., has more knowledge or computational power) than what is defined in the threat model. This subsection reminds the terminology commonly used in security literature (evans_pragmatic_2018).
@@
Secure protocols typically guarantee data privacy against a proportion of malicious/semi-honest agents. For instance, Clarion (eskandarian_clarion_2022) is secure as long as one out of $M$ servers (involved in the protocol) is honest. On the other hand, RPM (lu_rpm_2023) is secure as long as there is a majority of honest servers involved in the protocol.
@@
While most secure protocols restrict their threat model to the agents actively involved in the protocol, anonymous communication protocols can also be the target of traffic analysis attacks (murdoch_low-cost_2005). Some works then include in their threat model a “global adversary” able to capture the network traffic of the agents. A “global adversary” is typically a state or an internet service provider.
@@ 3.2. Shuffler definition
Our starting point is the definition of a shuffler $\mathcal{S}$. Based on papers studying shuffling-based DP (cheu_distributed_2019; cheu_differential_2021), we propose the following naive definition:
@@
Figure 1. Abstract shuffler design: (1) The senders pre-process their messages using $\mathcal{P}$, (2) They send their preprocessed messages to the mixer $\mathcal{M}$, (3) The mixer outputs a permutation of the messages, and (4) an analyser can use the shuffled messages in a privacy-preserving computation.: Schema of an abstract shuffler
@@
This informal definition fits what the literature (kairouz_advances_2021) implicitly considers as a “secure shuffler” (also referred to as “oblivious shuffler” in (bittau_prochlo_2017; gordon_spreading_2022)). However, this definition is imprecise. Thus, we propose to formalize this notion using four new security properties to characterize a shuffler $\mathcal{S}$: Anonymity, Correctness, Sender disruption resistance, Server disruption resistance.
@@ 3.3.1. Anonymity
This is the minimum requirement for shuffling-based private computations. More precisely, one is usually interested in sender anonymity because only the origin of the messages must remain anonymous. It is even the core concern of shuffler application paper that give more importance to the concept of “anonymous messaging” than to shuffling itself (ishai_cryptography_2006; cheu_distributed_2019; ghazi_pure_2020).
@@
While the first anonymization techniques (chaum1981untraceable; chaum_dining_1988) appeared in the 1980s, the first anonymity formalization paper (feigenbaum_model_2007) was only published in 2007. Thus, most anonymization systems (up until now) have been built upon informal anonymity definitions. For example, even recent constructions such as Clarion (eskandarian_clarion_2022) prove the security of their MPC protocol, but do not prove the anonymity itself; (eskandarian_clarion_2022) does not even cite any formalization framework.
pfitzmann_terminology_2010 proposed a recognized terminology for all the words related to anonymity (lu_survey_2019). These definitions are informal but are a commonly accepted starting point. They give the following definition for sender anonymity:
@@
Finding a formal definition for anonymity is a complex problem because anonymity is context-dependent. Several papers (backes_anoa_2013; gelernter_limits_2013; kuhn_privacy_2019) proposed formal anonymity definitions, but they remain imperfect. For example, some of them exclude some de-anonymization attacks (e.g., traffic analysis attacks in (backes_anoa_2013)), which creates a blind spot in the theory that struggles to fit the reality of anonymity. Furthermore, lu_survey_2019 surveyed these works on anonymity formalizations and concluded that the formalization proposed “is still a theoretic approach and lacks a practical verification.” Hence, anonymity formalization is still an open problem for the security community that is still looking for a standard and general solution.
@@
This definition means that an adversary can distinguish honest senders only with negligible probability. Suppose all the messages are unique, the probability of $m$ being sent by the honest party $P$ (i.e., $\Pr[m\text{ sent by }P|\mathcal{V}(A)]$) has two possible values: 0 if the adversary controls the sender of $m$, and $\frac{1}{\lvert\mathcal{H}\rvert}$, otherwise. If the messages are not all unique, we can have $q$ honest senders who sent the same message $m$; the probability of a (honest) $P$ being the sender of $m$ is $\frac{q}{\lvert\mathcal{H}\rvert}$.
@@
Our survey focuses on sender anonymity as it is the form of anonymity required by secure shufflers. However, some anonymous communication systems provide other forms of anonymity such as one-to-one anonymous communications (van_den_hooff_vuvuzela_2015; eskandarian_express_2021) (i.e., the two communicating agents know the identity of their correspondent, but there is a guarantee that no third party knows these agents are communicating). This is useful in a number of applications but incompatible with sender anonymity since the receiver knows the sender’s identity. Hence, such systems cannot be used as secure shufflers. Among all anonymous communication systems, only “anonymous broadcasters” satisfy the secure shuffling needs.
Over the years, several works including (edman2009anonymity; sasy_sok_2023; shirali_survey_2022; shirazi2018survey) have surveyed anonymous communication systems. As sasy_sok_2023 is the most recent and complete survey, we will focus our discussion on it, but our remarks hold for all surveys on anonymous communications. While both (sasy_sok_2023) and our work gravitates around the notion of anonymity, they provide different perspectives because they do not focus on the same functionality. On the one hand, they focus on anonymous communications. On the other hand, we focus on secure shuffling.
Since we focus on different functionalities, our work relies on a different set of definitions. For example, the concept of correctness (defined in the next subsubsection) is not studied in (sasy_sok_2023). Indeed, correctness is an important notion in privacy-preserving computations as we want to avoid a bias in the output results. The anonymous communications literature is much more interested in the notion of censorship resistance (which is implied by correctness but is not equivalent).
This scope difference has a visible consequence: while some solutions are covered in both works, some solutions are only present in (sasy_sok_2023) (e.g., one-to-one anonymous communications), and others are only present in our work (e.g., TEE-based shufflers).
A core contribution of our work is to identify secure protocols able to implement a secure shuffler, even though these protocols were not initially designed for it. While some anonymous communications systems can implement this functionality, many other secure protocols provide the sender anonymity necessary to secure shuffling. For example, Section 4 highlights ZKP protocols initially used for electronic voting (adida_helios_2008) and TEE solutions imagined for secure contact discovery (ngai_distributed_2024).